OpenStudy (anonymous):

Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)?

OpenStudy (anonymous):

x + 2y = 15 x + y = 15 x + y = 8 x − y = 2